# Safe In-Context Reinforcement Learning

### v1

@@ 5 Related Works
Learning to improve on a new MDP by interacting with it and without parameter updates is first studied in meta RL (Duan et al., 2016; Wang et al., 2016). See Beck et al. (2025) for a detailed survey of meta RL. In general, the demonstrated zero parameter update generalization is weak in meta RL literature. Most of them rely on task identification, i.e., identifying pretraining tasks that are similar to the evaluation task and acting as if the evaluation task were the pretraining tasks. Laskin et al. (2023) coined the word ICRL and demonstrated generalization to evaluation tasks that are far away from the pretraining tasks, which spurs increasing attention in ICRL. Variations of algorithm distillation have been proposed for efficient in-context reinforcement learning, including the Decision Pretrained Transformer (Lee et al., 2023) and Algorithm Distillation with Noise (Zisman et al., 2023). Both approaches aim to efficiently train transformer models for in-context learning using optimal policies. See Moeini et al. (2025) for a detailed survey of ICRL. The main novelty of this work is to first study safe ICRL within the constrained MDP framework.
Safe RL.
@@
Safe RL methods commonly adopt CMDP formulations to enforce compliance with safety constraints during exploration and policy optimization (Garcıa & Fernández, 2015; Gu et al., 2022; Wachi et al., 2024). Constrained Policy Optimization (CPO) serves as a foundational algorithm in this area, effectively balancing performance rewards with safety requirements (Achiam et al., 2017; Wachi & Sui, 2020). Subsequent extensions, such as Constrained Update Projection and Constrained Reward Policy Optimization, have built upon CPO to enhance theoretical guarantees and practical applicability in safe RL scenarios (Xu et al., 2021; Yang et al., 2022). Shielding with function encoders and conformal prediction has been proposed to handle unseen OOD environments (Kwon et al., 2025). However, this work emphasizes runtime adaptation rather than learning algorithms through contextual information. Reward-Constrained Policy Optimization (Tessler et al., 2018) is the closest work to EPPO, but it enforces safety only in expectation rather than at the episode level. From the offline RL perspective, the Constrained Decision Transformer (Liu et al., 2023) leverages the transformer architecture for safe RL. However, existing approaches enforce safety constraints only when the evaluation task closely matches the pretraining tasks. In contrast, our safe ICRL framework verifies safety constraints even on evaluation tasks that differ substantially from the pretraining distribution.
Safe Meta RL.
@@
Safe meta RL has emerged to enable fast adaptation to new tasks while enforcing safety constraints. Earlier work (Luo et al., 2021) uses a three-phase approach that meta-learns a safety critic from offline data across multiple environments, adapts it to a new environment with a smaller offline dataset, and employs the adapted safety critic with a recovery policy to ensure safe learning while minimizing constraint violations. Khattar et al. (2023) provides task-averaged regret bounds for rewards and constraints via gradient-based meta-learning. More recently, Guan et al. (2024) proposes a cost-aware context encoder that uses supervised cost relabeling and contrastive learning to infer tasks based on safety constraints. To our knowledge, all previous safe meta RL works require parameter updates to achieve safe adaptation in evaluation tasks. By contrast, our work on safe ICRL enables safe adaptation to evaluation tasks without parameter updates.
@@ 6 Conclusion
This work pioneers the study of safe ICRL, examining both reinforcement pretraining and supervised pretraining. We develope a safe supervised pretraining appraoch based on the constrained decision transformer, showcasing its capability for OOD generalization while revealing its dependence on dataset size. We introduce EPPO, a safe reinforcement pretraining approach via exact penalty policy optimization to enforce strict cost constraints within each episode. Theoretically, we prove that EPPO’s fixed point satisfies strict safety constraints while maximizing rewards. We conduct a thorough empirical study of the proposed approaches in challenging benchmarks that provably demand OOD generalization (Proposition 1). Such demand for OOD generalization is missing in many benchmarks used by previous works (Laskin et al., 2023; Zisman et al., 2023; Son et al., 2025) Through those challenging benchmarks, we confirm that our pretrained safe ICRL agents adapt to not only unseen but also OOD evaluation tasks while respecting the safety constraints. We envision this work guiding the development of robust, generalizable, and safe RL algorithms for real-world applications.
@@ Appendix A Algorithm
1:Input: discount factor $\gamma$, cost threshold $\delta$, number of training steps $T_{\max}$, batch size NN, env time limit $t_{\max}$, episodes-per-history range $K_{\min},K_{\max}]$, CTG range $\mathrm{CTG}_{\min},\mathrm{CTG}_{\max}]$, environment distribution $\mathcal{E}$
@@
4:  $K\leftarrow\texttt{rnd}(K_{\min},K_{\max})$,  $\mathrm{CTG}\leftarrow\texttt{rnd}(\mathrm{CTG}_{\min},\mathrm{CTG}_{\max})$,  $t\leftarrow 0$
$H\leftarrow[\;]$ // list of episodes
@@
35:   $d\lambda\leftarrow d\lambda+C_{e}^{\max}$
@@ B.1 Proof of Theorem 1
Let $J(\pi)\coloneqq{\mathbb{E}_{\pi}\quantity[\sum_{k=1}^{K}G(\tau_{k})]}$ in this proof. We prove the theorem in both directions. Also let $g_{k}^{+}(\pi)\coloneqq\max\{0,g_{k}(\pi)\}$, and $s(\pi)\coloneqq\max_{k}g_{k}(\pi)$.
@@
$L_{\Sigma}\big(\pi,\ \|\lambda^{\star}\|_{\infty}\big)\;=\;J(\pi)-\|\lambda^{\star}\|_{\infty}\sum_{i}g_{i}^{+}(\pi)\;\leq\;J(\pi)-\sum_{i}\lambda_{i}^{\star}g_{i}(\pi)\;\leq\;p^{\star}.$ (9)
@@
Together with (9), this shows
@@
If $s(\pi_{t+1})>0$, then $\lambda_{t+1}=[\lambda_{t}+\eta s(\pi_{t+1})]_{+}\geq\lambda_{t}$. If constraints are strict ($s(\pi_{t+1})<0$), then $\lambda_{t+1}\leq\lambda_{t}$. Under persistent violation, $\lambda_{t}$ eventually exceeds $\|\lambda^{*}\|_{\infty}$, by the exact-penalty bound (9), any maximizer of $L_{\Sigma}(\cdot,\lambda_{t})$ is then feasible/optimal and $s(\pi_{t+1})\leq 0$, so $\lambda$ stabilizes. ∎
@@ B.2 Proof of Proposition 1
$\mathbb{P}_{\text{train}}((i,j))=\frac{e^{-\alpha((i-i_{c})^{2}+(j-j_{c})^{2})}}{Z_{-\alpha}}\quad\text{and}\quad\mathbb{P}_{\text{test}}((i,j))=\frac{e^{\alpha((i-i_{c})^{2}+(j-j_{c})^{2})}}{Z_{\alpha}},$
@@
$D_{\mathrm{KL}}\left(\mathbb{P}_{\text{train }}\|\mathbb{P}_{\text{test }}\right)$ $=\sum_{(i,j)\in\mathcal{O}}\mathbb{P}_{\text{train }}((i,j))\log\frac{\mathbb{P}_{\text{train }}((i,j))}{\mathbb{P}_{\text{test }}((i,j))}$ (12)
$=\sum_{(i,j)\in\mathcal{O}}\mathbb{P}_{\text{train }}((i,j))\log\frac{Z_{\alpha}e^{-\alpha((i-i_{c})^{2}+(j-j_{c})^{2})}}{Z_{-\alpha}e^{\alpha((i-i_{c})^{2}+(j-j_{c})^{2})}}$ (13)
@@ Appendix C Training Details
Reinforcement Pretraining.
@@
In practice, since the environments are generated randomly, it is possible that an environment doesn’t have a feasible solution, causing the agent to inevitably violate the cost threshold. This results in $\lambda$ growing too large in an attempt to mitigate these cases. To avoid this, we suggest capping $\lambda$ to a reasonably large value. Ignoring abnormally high-cost episodes is another solution worth exploring.
Parameter SafeDarkRoom SafeDarkMujoco-Point SafeDarkMujoco-Car
$K_{\min}$, $K_{\max}$ 50, 50 20, 20 20, 20
Episode time limit $t_{\max}$ 30 75 75
Replay buffer capacity 100,000 100,000 100,000
Embedding Dim 64 64 64
Hidden Dim 64 64 64
Num Layers 4 4 4
Num Heads 8 8 8
Seq Len 1500 1500 1500
Attention Dropout 0 0 0
Residual Dropout 0 0 0
Embedding Dropout 5 5 5
Learning Rate 3e-4 3e-4 3e-4
Betas (0.9, 0.99) (0.9, 0.99) (0.9, 0.99)
Clip Grad 1.0 1.0 1.0
Batch Size 32 32 32
Num Updates 30k 10k 10k
Optimizer Adam Adam Adam
Table 1: Parameters for Safe Reinforcement Pretraining
Supervised Pretraining.
@@
In supervised pretraining, we collect a dataset $\mathcal{D}=\quantity{\Xi_{i}}$ comprising multiple trajectories, where each trajectory $\Xi_{i}\doteq(\tau_{1},\tau_{2},\dots,\tau_{K})$ represents a sequence of episodes generated by running existing safe RL algorithms on various CMDPs. Each episode $\tau_{k}$ in a trajectory $\Xi_{i}$ comprises states, actions, rewards, and costs, with the episode return $G(\tau_{k})\doteq\sum_{t=1}^{T}R_{t}$ and cost-to-go $G_{c,t}(\tau_{k})\doteq\sum_{i=t+1}^{T}C_{i}$, expected to increase and decrease, respectively, with $k$ as the RL algorithm learns. The policy $\pi_{\theta}$ is trained autoregressively using an imitation learning loss to distill the behavior of RL algorithms demonstrated in the dataset, following the concept of algorithm distillation (Laskin et al., 2023). For discrete action spaces (e.g., SafeDarkRoom), we use a cross-entropy loss, defined as:
$\mathcal{L}_{\mathrm{CE}}(\theta)=\mathbb{E}_{\Xi_{i}\sim\mathcal{D}}\left[-\log\pi_{\theta}(A_{t}^{k}\mid S_{t}^{k},H_{t}^{k},G_{t}(\tau_{k}),G_{c,t}(\tau_{k}))\right],$ (16)
@@
where $\pi_{\theta}$ is the categorical distribution over discrete actions. For continuous action spaces (e.g., SafeDarkMujoco), we use an L2 loss, where the transformer outputs an action mean $\mu_{t}=\mu_{\theta}(S_{t}^{k},H_{t}^{k},G_{t}(\tau_{k}),G_{c,t}(\tau_{k}))$, minimizing:
$\mathcal{L}_{\mathrm{L2}}(\theta)=\mathbb{E}_{\Xi_{i}\sim\mathcal{D}}\left[\left\|A_{t}^{k}-\mu_{\theta}(S_{t}^{k},H_{t}^{k},G_{t}(\tau_{k}),G_{c,t}(\tau_{k}))\right\|^{2}\right].$ (17)
These losses extend the imitation learning loss from Equation (5) by incorporating RTG and CTG conditioning, aligning with our goal of distilling RL algorithms into the policy’s forward pass while optimizing for safety constraints.
@@
Dataset Collection. For supervised pretraining, we collect learning trajectories using reinforcement learning (RL) algorithms. As our base RL algorithm, we employ PPO-Lagrangian Schulman et al. (2017); Ray et al. (2019), which is designed to maximize rewards while enforcing safety constraints. These trajectories capture behaviors that learn to avoid obstacles.
To introduce variation in the learned behaviors, we vary the cost limits in PPO-Lag across multiple settings. For the SafeDarkRoom environment, we use three cost limits: 0, 2.5, and 5.0. For the SafeDarkMujoco environment, we similarly use three cost limits: 0, 5, and 10. For each cost limit, we collect 50,000 steps of learning history in both environments.
@@
We report the hyperparameters used for reinforcement pretraining (Table 1) and supervised pretraining (Table 2). Reinforcement pretraining adopts the AMAGO framework (Grigsby et al., 2024a) integrated with our novel EPPO method. Supervised pretraining employs a constrained decision transformer as the backbone (Liu et al., 2023).
Parameter SafeDarkRoom SafeDarkMujoco-Point SafeDarkMujoco-Car
Embedding Dim 64 64 64
Hidden Dim 512 256 256
Num Layers 8 8 8
Num Heads 8 8 8
Seq Len 100 300 200
Attention Dropout 0.5 0.5 0.5
Residual Dropout 0.1 0.1 0.1
Embedding Dropout 0.3 0.3 0.3
Learning Rate 3e-4 3e-4 3e-4
Betas (0.9, 0.99) (0.9, 0.99) (0.9, 0.99)
Clip Grad 1.0 1.0 1.0
Batch Size 512 128 128
Num Updates 300k 500k 500k
Optimizer Adam Adam Adam
Table 2: Parameters for Safe Supervised Pretraining
Appendix D Variation of Supervised Pretraining
Figure 4: Performance Comparison of Algorithm Distillation (AD) style supervised pretraining and Algorithm Distillation with Noise (AD-EPS) style supervised pretraining in SafeDarkRoom. AD-EPS fails to generalize in OOD environments.
In this section, we compare the performance of Algorithm Distillation (AD) (Laskin et al., 2023) and Algorithm Distillation with Noise (AD-EPS) (Zisman et al., 2023) in SafeDarkRoom. AD-EPS is designed to learn in-context RL algorithms using datasets generated from a perturbed optimal policy. This approach allows efficient generation of learning trajectories from a single optimal policy. However, our results in complex environments with cost signals reveal that AD-EPS relies on artificial trajectories. AD-EPS fails to learn effective in-context RL algorithms when relying solely on perturbed optimal policies, as the perturbations do not account for obstacle avoidance in SafeDarkRoom environment and instead introduce random action variations.

### v2

@@ 6 Related Works
ICRL. Learning to improve on a new MDP by interacting with it without parameter updates was first studied in meta-RL (Duan et al., 2016; Wang et al., 2016). See Beck et al. (2025) for a detailed survey of meta-RL. In general, zero-parameter-update generalization in meta-RL is limited, as most methods rely on task identification—matching evaluation tasks to similar pretraining tasks, i.e., identifying pretraining tasks that are similar to the evaluation task and acting as if the evaluation task were the pretraining tasks. Laskin et al. (2023) coined the term ICRL and demonstrated strong generalization to tasks far from the pretraining distribution, spurring growing interest in the area.
Variations of algorithm distillation have been proposed for efficient in-context reinforcement learning, including the Decision Pretrained Transformer (Lee et al., 2023) and Algorithm Distillation with Noise (Zisman et al., 2023). Both approaches aim to efficiently train transformer models for in-context learning using optimal policies. LLMs have been shown to exhibit ICRL as an emergent capability, enabling them to keep improving on tasks when previous attempts are given along with a reward in the context (Song et al., 2025). See Moeini et al. (2025) for a detailed survey of ICRL. However, none of these works have investigated safety within the context of ICRL.
@@
Safe RL. Safe RL methods commonly adopt CMDP formulations to enforce compliance with safety constraints during exploration and policy optimization (García & Fernández, 2015; Gu et al., 2022; Wachi et al., 2024). Constrained Policy Optimization (CPO) serves as a foundational algorithm in this area, balancing rewards with safety (Achiam et al., 2017; Wachi & Sui, 2020). Building on CPO, subsequent works enhance theoretical guarantees and scalability through primal-dual formulations and efficient projection-based updates (Tessler et al., 2019; Xu et al., 2021; Yang et al., 2022). Shielding with function encoders and conformal prediction has been proposed to handle unseen OOD environments (Kwon et al., 2025). However, this work emphasizes runtime adaptation rather than learning algorithms through interaction. From the offline RL perspective, Constrained Decision Transformer (Liu et al., 2023) uses Transformer architecture for safe RL. However, existing approaches enforce safety constraints only when the evaluation task closely matches the pretraining tasks. In contrast, our safe ICRL framework satisfies safety constraints even on evaluation tasks that differ substantially from the pretraining distribution.
@@
Safe Meta RL. Safe meta RL enables fast adaptation to new tasks, enforcing safety constraints. For example, Luo et al. (2021) propose a three-phase strategy that meta-learns a safety critic across environments, adapts it to new tasks using limited offline data, and use it with a recovery policy to reduce constraint violations during learning. Khattar et al. (2023) provides task-averaged regret bounds for rewards and constraints via gradient-based meta-learning. More recently, Guan et al. (2024) proposes a cost-aware context encoder that uses supervised cost relabeling and contrastive learning to infer tasks based on safety constraints. Xu & Zhu (2025) proposes a safe meta-RL algorithm that integrates a closed-form, one-step safe policy adaptation mechanism to provide anytime safety guarantees. Unlike safe meta-RL methods, which rely on parameter updates to achieve safe adaptation during evaluation, safe ICRL enables adaptation without test-time parameter updates.
@@ 7 Conclusion
This work pioneers the study of safe ICRL, where agents must adapt to new and potentially OOD tasks without parameter updates while adhering to safety constraints. We show that existing approaches, including algorithm distillation for safe RL, and safe meta-RL baselines, are limited in this setting, as they lack mechanisms for regulating safety through in-context interaction histories and for adjusting to cost budgets during adaptation.
To address these limitations, we introduced SCARED, a reinforcement pretraining approach for safe in-context adaptation under the CMDP framework. SCARED enables agents to balance reward maximization and cost minimization during adaptation by regulating episode-level safety constraints and adjusting behavior in response to different cost budgets. Theoretically, we prove that fixed points of SCARED optimization satisfy strict safety constraints while maximizing rewards. To rigorously test OOD generalization, we introduce challenging benchmarks for safe ICRL (Proposition 1), addressing a gap in prior evaluations (Laskin et al., 2023; Zisman et al., 2023; Son et al., 2025). Through these challenging benchmarks, we confirm that our pretrained safe ICRL agents adapt to not only unseen ID but also OOD evaluation tasks while respecting safety constraints. We envision this work guiding the development of robust, generalizable, and safety-aware in-context learning methods for real-world applications.
@@ Acknowledgments
Broader Impacts Statement
This work introduces the first study of safety in ICRL, a paradigm in which agents adapt to new tasks at deployment time without parameter updates. By providing a framework for enforcing user-specified safety budgets during test-time adaptation, this work is relevant to real-world applications such as robotics and embodied AI, where retraining is impractical and unsafe exploration can have real consequences.
The proposed framework does not eliminate risk, particularly in environments with unmodeled hazards or adversarial conditions, and should be complemented by human oversight and additional safety mechanisms. We release our benchmarks and evaluation protocols to support transparent and reproducible research on safe adaptation under distribution shift.
@@ Appendix A Algorithm
Algorithm 1 SCARED Implemented with DDPG
1:Input: discount factor $\gamma$, cost budget $\delta$, number of training steps $T_{\max}$, batch size NN, env time limit $t_{\max}$, episodes-per-history range $K_{\min},K_{\max}]$, CTG range $\mathrm{CTG}_{\min},\mathrm{CTG}_{\max}]$, environment distribution $\mathcal{E}$
@@
4:  $K\leftarrow\texttt{rnd}(K_{\min},K_{\max})$,  $\mathrm{CTG}\leftarrow\texttt{rnd}(\mathrm{CTG}_{\min},\mathrm{CTG}_{\max})$,  $t\leftarrow 0$ $H\leftarrow[\;]$ // list of episodes
@@
35:   $d\lambda\leftarrow d\lambda+(C_{e}^{\max}-\delta)$
@@ B.1 Proof of Theorem 1
Let $J(\pi)\coloneqq{\mathbb{E}_{\pi}\hskip-1.0pt\quantity[\sum_{k=1}^{K}\hskip-1.0ptG(\tau_{k})]}$ in this proof. We prove the theorem in both directions. Also let $g_{k}^{+}(\pi)\coloneqq\max\{0,g_{k}(\pi)\}$, and $s(\pi)\coloneqq\max_{k}g_{k}(\pi)$.
@@
$L_{\Sigma}\big(\pi,\ \|\lambda^{\star}\|_{\infty}\big)\;$ $=\;J(\pi)-\|\lambda^{\star}\|_{\infty}\sum_{i}g_{i}^{+}(\pi)$
$\leq\;J(\pi)-\sum_{i}\lambda_{i}^{\star}g_{i}(\pi)$
$\leq\;p^{\star}.$ (8)
@@
Together with (B.1), this shows
@@
If $s(\pi_{t+1})>0$, then $\lambda_{t+1}=[\lambda_{t}+\eta s(\pi_{t+1})]_{+}\geq\lambda_{t}$. If constraints are strict ($s(\pi_{t+1})<0$), then $\lambda_{t+1}\leq\lambda_{t}$. Under persistent violation, $\lambda_{t}$ eventually exceeds $\|\lambda^{*}\|_{\infty}$, by the exact-penalty bound (B.1), any maximizer of $L_{\Sigma}(\cdot,\lambda_{t})$ is then feasible/optimal and $s(\pi_{t+1})\leq 0$, so $\lambda$ stabilizes. ∎
Note: The full convergence proof to an exact fixed point is beyond the scope of this work.
@@ B.2 Proof of Proposition 1
[…]
$\mathbb{P}_{\text{test}}((i,j))=\frac{e^{\alpha((i-i_{c})^{2}+(j-j_{c})^{2})}}{Z_{\alpha}}$
@@
$D_{\mathrm{KL}}\left(\mathbb{P}_{\text{train }}\|\mathbb{P}_{\text{test }}\right)$
$=\sum_{(i,j)\in\mathcal{O}}\mathbb{P}_{\text{train }}((i,j))\log\frac{\mathbb{P}_{\text{train }}((i,j))}{\mathbb{P}_{\text{test }}((i,j))}$
$=\sum_{(i,j)\in\mathcal{O}}\mathbb{P}_{\text{train }}((i,j))\log\frac{Z_{\alpha}e^{-\alpha((i-i_{c})^{2}+(j-j_{c})^{2})}}{Z_{-\alpha}e^{\alpha((i-i_{c})^{2}+(j-j_{c})^{2})}}$
@@ Appendix C Training Details
For SafeVelocity, we use the Ant and HalfCheetah environments from SafetyGym (Ji et al., 2023), but sample a different target velocity for each trajectory to convert them into safe ICRL benchmarks. Other aspects of the environments remain unchanged.
SCARED.
@@
Parameter SafeDarkRoom SafeDarkMujoco & SafeVelocity
$K_{\min}$, $K_{\max}$ 50, 50 20, 20
Episode time limit $t_{\max}$ 30 75
Replay buffer capacity 100,000 100,000
Embedding Dim 64 64
Hidden Dim 64 64
Num Layers 4 4
Num Heads 8 8
Seq Len 1500 1500
Attention Dropout 0 0
Residual Dropout 0 0
Embedding Dropout 5 5
Learning Rate 3e-4 3e-4
Betas (0.9, 0.99) (0.9, 0.99)
Clip Grad 1.0 1.0
Batch Size 32 32
Num Updates 30k 10k
Optimizer Adam Adam
Table 1: Parameters for SCARED
SafeMeta & MAML with penalty
We use the original implementations by (Xu & Zhu, 2025) with 3 hidden layers of (64, 512, 64) units for policy, value, and cost networks, totaling 205,510 parameters, comparable to SCARED. We train for 15k meta-iterations, where each iteration samples 20 tasks from the training distribution. During meta-testing, both algorithms perform $K$ gradient updates to adapt to new tasks with unseen cost-to-go values. Note that while some meta-RL methods use frame stacking or recurrent architectures, their hidden states reset between episodes, making them fundamentally episodic—they do not carry information across episode boundaries as ICRL methods do.
Parameter SafeDarkRoom SafeDarkMujoco & SafeVelocity
$K_{\min}$, $K_{\max}$ 50, 50 20, 20
Episode time limit $t_{\max}$ 30 75
Hidden layers (64, 512, 64) (64, 512, 64)
Min/Max batch size 500/1500 1500/1500
Policy learning rate 1e-3 1e-3
Value/Cost learning rate 3e-2/1e-1 3e-2/1e-1
Discount factor $\gamma$ 0.99 0.99
GAE parameter $\tau$ 0.95 0.95
Max KL divergence 1e-3 1e-3
Lagrangian weight $\lambda$ 1.0 1.0
Num meta-training iterations 15k 15k
Policy optimizer Adam Adam
Value/Cost optimizer L-BFGS L-BFGS
Table 2: Parameters for SafeMeta and MAML with penalty
Safe Algorithm Distillation.
@@
In safe algorithm distillation, we collect a dataset $\mathcal{D}=\quantity{\Xi_{i}}$ comprising multiple trajectories, where each trajectory $\Xi_{i}\doteq(\tau_{1},\tau_{2},\dots,\tau_{K})$ represents a sequence of episodes generated by running existing safe RL algorithms on various CMDPs. Each episode $\tau_{k}$ in a trajectory $\Xi_{i}$ comprises states, actions, rewards, and costs, with the episode return $G(\tau_{k})\doteq\sum_{t=1}^{T}R_{t}$ and cost-to-go $G_{c,t}(\tau_{k})\doteq\sum_{i=t+1}^{T}C_{i}$, expected to increase and decrease, respectively, with $k$ as the RL algorithm learns.
We train the policy $\pi_{\theta}$ autoregressively using safe algorithm distillation to distill the behavior of safe RL algorithms present in the dataset, following Laskin et al. (2023). Concretely, each training example consists of a trajectory segment $\Xi_{i}$ containing state and action pairs along with RTG and CTG conditioning. Let the policy take as input $(S_{t}^{k},H_{t}^{k},G_{t}(\tau_{k}),G_{c,t}(\tau_{k}))$ and predict the action at time $t$ for episode $k$.
Discrete action spaces. For environments with discrete actions (e.g., SafeDarkRoom), the model outputs categorical logits $z_{\theta}(\cdot)$, generating a categorical distribution $\pi_{\theta}(\cdot\mid\cdot)=\mathrm{Softmax}(z_{\theta}(\cdot))$. We train with the standard cross-entropy loss:
$\mathcal{L}_{\mathrm{disc}}(\theta)$ $=\mathbb{E}_{\Xi_{i}\sim\mathcal{D}}\left[-\log\pi_{\theta}\!\left(A_{t}^{k}\mid S_{t}^{k},H_{t}^{k},G_{t}(\tau_{k}),G_{c,t}(\tau_{k})\right)\right].$ (13)
@@
Continuous action spaces. For environments with continuous actions (e.g., SafeDarkMujoco), the model outputs a conditional distribution over actions parameterized by a mean $\mu_{\theta}(\cdot)$. We supervise the predicted mean using an $\ell_{2}$ loss:
$\mathcal{L}_{\mathrm{cont}}(\theta)$ $=\mathbb{E}_{\Xi_{i}\sim\mathcal{D}}\left[\left\|A_{t}^{k}-\mu_{\theta}\!\left(S_{t}^{k},H_{t}^{k},G_{t}(\tau_{k}),G_{c,t}(\tau_{k})\right)\right\|_{2}^{2}\right].$ (14)
These objectives enables the transformer to distill constraint-aware goal seeking behaviors into its forward pass.
Dataset Collection.
@@
For safe AD, we collect learning trajectories using safe RL algorithms. As our base safe RL algorithm, we employ PPO-Lagrangian (Schulman et al., 2017; Ray et al., 2019), which is designed to maximize rewards while enforcing safety constraints. These trajectories capture behaviors that learn to avoid obstacles.
To introduce variation in the learned behaviors, we vary the cost limits in PPO-Lag across multiple settings. For the SafeDarkRoom environment, we use three cost limits: 0, 2.5, and 5.0. For each cost limit, we collect 50,000 steps of learning history. For the SafeDarkMujoco and SafetyVelocity environments, prior offline RL benchmarks and constrained decision transformer methods typically use datasets on the order of 1122 million environment steps (Fu et al., 2020; Liu et al., 2023). In our experiments, we intentionally use larger datasets of 2255 million steps. This difference from the usual convention is motivated by the need to collect diverse cost-related experiences in continuous control settings, where meaningful safety-constrained behavior emerges only after long training horizons.
In our data collection, we set a single cost limit of 0 throughout the full learning history. This setup intentionally includes trajectories generating a various range of cost violations, from high-cost exploration phases to near-zero-cost behavior as training progresses. Continuous control tasks are substantially more difficult to solve than discrete ones, and learning meaningful policies under safety constraints typically requires more interaction data.
Thus, this design poses a challenge for safe supervised in-context RL: conditioning on both return-to-go and cost-to-go induces a large combinatorial space of reward–cost trade-offs that must be represented in the dataset. As a result, the required dataset size grows substantially.
@@
We report the hyperparameters used for SCARED (Table 1) and safe AD (Table 3). SCARED adopts the AMAGO framework (Grigsby et al., 2024a) integrated with our SCARED method. Safe AD employs a constrained decision transformer as the backbone (Liu et al., 2023).
Parameter SafeDarkRoom SafeDarkMujoco & SafeVelocity
Embedding Dim 64 512
Hidden Dim 512 256
Num Layers 8 8
Num Heads 8 8
Seq Len 100 200
Attention Dropout 0.5 0.5
Residual Dropout 0.1 0.1
Embedding Dropout 0.3 0.3
Learning Rate 3e-4 3e-4
Betas (0.9, 0.99) (0.9, 0.99)
Clip Grad 1.0 1.0
Batch Size 512 128
Num Updates 300k 500k
Optimizer Adam Adam
Table 3: Parameters for Safe Algorithm Distillation
Appendix D Variation of Safe Algorithm Distillation
Figure 4: Performance comparison between algorithm distillation and algorithm distillation with noise in SafeDarkRoom. AD-EPS fails to generalize to out-of-distribution (OOD) environments.
In this section, we compare safe algorithm distillation (AD) (Laskin et al., 2023) and safe algorithm distillation with noise (AD-EPS) (Zisman et al., 2023) in SafeDarkRoom environment. AD-EPS aims to learn in-context RL algorithms by training on datasets generated from a single optimal policy with injected action noise, enabling efficient collection of learning trajectories.
However, in environments with safety constraints and explicit cost signals, we find that AD-EPS struggles to generalize. In SafeDarkRoom, the perturbed trajectories produced by AD-EPS fail to capture meaningful obstacle-avoidance behavior, as the injected noise primarily induces random action variations rather than structured exploration. As a result, AD-EPS relies on artificial trajectories that do not reflect the true learning dynamics required for effective in-context RL, leading to poor generalization in out-of-distribution settings.